# Leveraging Large Language Models for Suicide Detection on Social Media with Limited Labels

### v2

@@ Abstract
large language models, text classification, limited labels, prompt engineering, suicide detection, social media analysis
@@ I Introduction
Deep learning techniques in natural language processing (NLP) have demonstrated significant potential in automating the identification of suicidal content in social media posts. These techniques, due to their ability to recognize subtle patterns in text data, are increasingly used to classify and detect posts indicating suicidal thoughts. The process typically involves three main tasks: collecting textual data from social media platforms, labeling the data, and building a deep learning-based classifier. The process of labeling is time-consuming and requires domain experts, which results in very limited annotated datasets. Additionally, social media posts may contain vague or implicit intent, necessitating a strong language understanding ability for accurate classification. These factors pose a challenge in developing an effective model for detecting suicide.
@@
In this study, we investigate the use of Large Language Models (LLMs) to classify the signs of suicide from user posts. We start by proposing a method for generating pseudo-labels for unlabeled data. We annotate user posts using LLMs (e.g., Qwen2-72B-Instruct [3]) using prompting. To minimize the noise in the labeling process, we fine-tune two more models (Llama3-8B [4] and DepRoBERTa [5]) on a small set of annotated data to filter out unreliable labels. The unlabeled data with pseudo-labels is combined with the small set of labeled data to form a new training set. We then fine-tune some more LLMs (e.g., Llama3-8B, Gemma2-9B) on the newly formed dataset, and evaluate the effectiveness of the models in suicide classification. Finally, we combine these fine-tuned models, together with prompting LLMs, to create an ensemble for a more robust and performant suicide detector. We apply our approach to the Suicide Ideation Detection on Social Media Challenge, a track in the IEEE Big Data 2024 Big Data Cup. Additionally, we provide analysis and discussion to gain insights into the results, such as how different LLMs affect the prompting performance, and choices of the loss function when fine-tuning. In summary, our contributions are:
@@ III-A Deep learning approaches for text classification
LLMs via prompt engineering have become a critical method for improving the capabilities of LLMs (e.g., LLaMA3 [4], GPT-4 [9]) across various downstream tasks [19]. In this approach, a prompt - that is, a task-specific instruction along with some examples (few-shot) and a question - can be directly input into LLMs to obtain a response for the question. This technique enhances the model’s effectiveness by utilizing in-context learning, without the need to modify the model parameters. To further enhance the reasoning ability of LLMs, Wang et al. introduced Self-Consistency, where a single LLM generates multiple responses to a given question and then uses majority voting to determine the final answer [20]. Wei et al. introduced Chain-of-Thought (CoT), a method that employs a series of intermediate reasoning steps to reach the final answer. In another line of work, Du et al. organized multiple LLMs in a debate setting, where each LLM independently generates an initial response, and then communicates to reach a consensus on a final answer [21]. Meanwhile, Pham et al. removed the token sampling step from LLMs, allowing LLMs to debate more effectively through the raw transformer output embeddings [22]. Overall, these techniques have been effective in improving the performance of LLMs through the use of prompts in downstream tasks. In our approach, we utilize LLMs with few-shot CoT prompting to generate pseudo-labels for unlabeled data, aiding in the creation of a substantial training set for the fine-tuning phase.
@@ III-B Large Language Models for Mental Health Detection
Recent research has expanded the use of large language models (LLMs) for detecting mental health issues. Lan et al. proposed combining LLMs and traditional classifiers to integrate medical knowledge-guided features, achieving both high accuracy and explainability in depression detection [23]. Poswiata et al. introduced fine-tuning RoBERTa [24] on depression-related posts to achieve an effective depression detector [5]. Yang et al. introduced MentalLLaMA based on LLaMA2 [25], for interpretable mental health analysis with instruction-following capability [26]. Arcan et al. provided a comprehensive evaluation of LLaMA2 and ChatGPT-4 for mental health tasks, revealing both the potential and limitations of LLM-based methods in this field [27]. There is another line of work in utilizing open-source LLMs to identify evidence in online posts that indicates the level of suicidal risks [28]. As for suicide detection, Li et al. proposed a new benchmark, consisting of fine-grained annotations of suicidal posts with BERT-based models [2]. Meanwhile, Qi et al. [29] presented another benchmark for Chinese social media datasets, assessing the performance of traditional models and Chinese LLMs such as Chinese-LLaMa-2-7B [30]. In this paper, we focus on using state-of-the-art LLMs under a limited labeled data setting to create an effective and reliable classifier for identifying suicidal content in English social media posts.
@@ IV-A Generating pseudo-labels for unlabeled data
When a new post is introduced, which replaces the placeholder, highlighted in orange (as shown in Fig. 2), the LLM is expected to generate a response in the same format as the exemplar, including three questions, three corresponding answers, and a compiled final answer, all within the dashed box (Fig. 2).
@@
Through error analysis of the annotation results, we observed that some posts involved writers who had previously attempted suicide but then expressed a desire to continue living and move forward. LLMs may misinterpret the overall context of these posts, leading to incorrect annotations as Attempts. To address this issue, we introduced an additional prompt specifically for posts initially labeled as Attempts at the initial round. The prompt template is illustrated in Fig. 3. This prompt checks whether the writer mentions moving on from suicidal thoughts or attempts, such as feeling happy to be alive or having found a reason to live. If the answer generated by LLM is “Yes,” we modify the original answers by flipping the last answer corresponding to Attempt to “No.” For example, if LLMs initially predict a post as {Yes, No, Yes} and the writer mentions moving on, we change the last “Yes” for Attempts to “No”. The refined answers then become {Yes, No, No}, and the post would be re-annotated as Ideation.
@@
We employ Qwen2-72B-Instruct  [^1] https://huggingface.co/Qwen/Qwen2-72B-Instruct to generate pseudo-label for the unlabeled posts, as it is one of the performant open-source models with an extended context window of up to $128K$ tokens [3]. We use greedy-decoding (i.e., choosing the highest probability output at each step of the sequence generation) and set the maximum length of the generated tokens to $1024$. We also experiment with other LLMs, such as LLaMA3 [4] and Mistral family models [37, 38], refer to Table VII and Section VI for further discussion.
@@
Combining models to generate training dataset. We use a combination of Qwen2-72B-Instruct, along with fine-tuned DepRoBERTa and Llama3-8B models on the $500$ labeled posts to generate pseudo-labels for unlabeled data, as depicted in Fig. 1(a), rightmost. Specifically, we only kept posts for which all three models predicted the same labels, and discarded the rest in case of disagreement. By doing so, we retain only high-confidence pseudo-labels, which allow us to reduce model-specific biases, resulting in a cleaner dataset for fine-tuning later on. Refer to Section VI for more discussion. After labeling $1{,}500$ posts, we keep only $902$ posts with pseudo-labels. These posts, combined with the $500$ labeled dataset, form the training set of $1{,}402$ posts. Statistics for the four annotated suicide risk labels from the combined dataset are outlined in Table III.
@@ IV-C Ensemble Model for a robust and performant classifier
Our ensemble model consisting of five individual models: Qwen2-72B-Instruct via prompting, and four fine-tuned large language models (LLMs) - Llama3-8B (with two variants differing in hyperparameters, denoted as Llama3-8B 1 and Llama3-8B 2), Llama3.1-8B, and Gemma2-9B, as shown in Fig. 1(c). From $500$ labeled data, we assign the ensemble weight $\mathrm{w}_{i}$ to the models, with a weight of $2$ for Qwen2-72B-Instruct, and a weight of $1$ for the other four models. This means that the prediction of Qwen2-72B-Instruct counts twice, whereas all other models count once. The final prediction $\mathrm{p}$ is determined through a weighted majority voting over the five models as illustrated in Eq. 4. This ensemble approach combines the strengths of each individual model to create a more performant and robust classifier, as demonstrated and discussed in Section V and Section VI.
@@ VI Analysis and Discussion
Figure 6: Correlation Between Model Predictions on 500 labeled Data. The predictions of individual models show a relatively low correlation with each other. This low correlation is a desirable attribute, as it typically leads to improved performance and robustness for the ensemble model.
@@
Correlation Between Model Predictions. As seen in Fig. 4, the ensemble model shows its robustness when tested on a new dataset. To gain further insight, we compared the correlation between the predictions of the models, as illustrated in Fig. 6. Creating an effective ensemble model requires not only high performance, but also diversity in the individual models. When these models have low correlation, it helps to reduce the risk of overfitting some parts of the data, leading to better generalization. The results show that the individual models produce diverse predictions with relatively low correlation, contributing to better generalization in the ensemble’s performance.
@@ VII Conclusion
However, our work has some limitations. The ensemble model, particularly Qwen2-72B-Instruct, requires a substantial amount of time for inference due to its large size, which poses challenges for real-time deployment. To mitigate this, smaller distillation models [44] could offer a more practical solution. Additionally, enhancing suicide detection using LLMs would benefit from better prompt engineering by domain experts. Breaking prompts into smaller, more detailed questions can help guide LLMs more effectively and interoperably. Another promising direction for future research is integrating visual data, such as images and videos commonly found in user posts, which may necessitate the use of multimodal LLMs.

### v3

@@ Abstract
large language models, text classification, limited labels, semi-supervised learning, prompt engineering, suicide detection, social media analysis
@@ I Introduction
Deep learning techniques in natural language processing (NLP) have shown promise in detecting suicidal content in social media posts by effectively recognizing subtle patterns in the text data. The process typically involves three main tasks: collecting textual data from social media platforms, labeling the data, and building a deep learning-based classifier. The process of labeling is time-consuming and requires domain experts, which results in very limited annotated datasets. Additionally, social media posts may contain vague or implicit intent, necessitating a strong language understanding ability for accurate classification. These factors pose a challenge in developing an effective model for detecting suicide.
@@
In this study, we investigate the use of Large Language Models (LLMs) to classify signs of suicide from user posts within a semi-supervised setup. We start by proposing a method for generating pseudo-labels for unlabeled data. We annotate user posts using LLMs (e.g., Qwen2-72B-Instruct [3]) using prompting. To minimize the noise in the labeling process, we fine-tune two more models (Llama3-8B [4] and DepRoBERTa [5]) on a small set of annotated data to filter out unreliable labels. The unlabeled data with pseudo-labels is combined with the small set of labeled data to form a new training set. We then fine-tune some more LLMs (e.g., Llama3-8B, Gemma2-9B) on the newly formed dataset, and evaluate the effectiveness of the models in suicide classification. Finally, we combine these fine-tuned models, together with prompting LLMs, to create an ensemble for a more robust and performant suicide detector. We apply our approach to the Suicide Ideation Detection on Social Media Challenge, a track in the IEEE Big Data 2024 Big Data Cup. Additionally, we provide analysis and discussion to gain insights into the results, such as how different LLMs affect the prompting performance, and choices of the loss function when fine-tuning. In summary, our contributions are:
@@ III-A Deep learning approaches for text classification
LLMs via prompt engineering have become a critical method for improving the capabilities of LLMs (e.g., LLaMA3 [4], GPT-4 [9]) across various downstream tasks [19]. In this approach, a prompt - that is, a task-specific instruction along with some examples (few-shot) and a question - can be directly input into LLMs to obtain a response for the question. This technique enhances the model’s effectiveness by utilizing in-context learning, without the need to modify the model parameters. To further enhance the reasoning ability of LLMs, Wang et al. [20] introduced Self-Consistency, where a single LLM generates multiple responses to a given question and then uses majority voting to determine the final answer. Wei et al. [21] introduced Chain-of-Thought (CoT), a method that employs a series of intermediate reasoning steps to reach the final answer. In another line of work, Du et al. [22] organized multiple LLMs in a debate setting, where each LLM independently generates an initial response, and then communicates to reach a consensus on a final answer. Meanwhile, Pham et al. [23] removed the token sampling step from LLMs, allowing LLMs to debate more effectively through the raw transformer output embeddings. Overall, these techniques have been effective in improving the performance of LLMs through the use of prompts in downstream tasks. In our approach, we utilize LLMs with few-shot CoT prompting to generate pseudo-labels for unlabeled data, aiding in the creation of a substantial training set for the fine-tuning phase.
@@ III-B Large Language Models for Mental Health Detection
Recent research has expanded the use of large language models (LLMs) for detecting mental health issues. Lan et al. proposed combining LLMs and traditional classifiers to integrate medical knowledge-guided features, achieving both high accuracy and explainability in depression detection [24]. Poswiata et al. introduced fine-tuning RoBERTa [25] on depression-related posts to achieve an effective depression detector [5]. Yang et al. [26] introduced MentalLLaMA based on LLaMA2 [27], for interpretable mental health analysis with instruction-following capability. Arcan et al. [28] provided a comprehensive evaluation of LLaMA2 and ChatGPT-4 for mental health tasks, revealing both the potential and limitations of LLM-based methods in this field. There is another line of work in utilizing open-source LLMs to identify evidence in online posts that indicates the level of suicidal risks [29]. As for suicide detection, Li et al. [2] proposed a new benchmark, consisting of fine-grained annotations of suicidal posts with BERT-based models. Meanwhile, Qi et al. [30] presented another benchmark for Chinese social media datasets, assessing the performance of traditional models and Chinese LLMs such as Chinese-LLaMa-2-7B [31]. In this paper, we focus on using state-of-the-art LLMs under a limited labeled data setting to create an effective and reliable classifier for identifying suicidal content in English social media posts.
@@ IV-A Generating pseudo-labels for unlabeled data
When a new post replaces the orange-highlighted placeholder (Fig. 2), the LLM is expected to generate a response in the same format as the exemplar, including three questions, three corresponding answers, and a compiled final answer, all within the dashed box (Fig. 2).
@@
Through error analysis of the annotation results, we observed that some posts involved writers who had previously attempted suicide but then expressed a desire to continue living and move forward. LLMs may misinterpret the overall context of these posts, leading to incorrect annotations as Attempts. To address this issue, we introduced an additional prompt specifically for posts initially labeled as Attempts at the initial round. The prompt template is illustrated in Fig. 3. This prompt checks whether the writer mentions moving on from suicidal thoughts or attempts, such as feeling happy to be alive or having found a reason to live. If the answer generated by LLM is “Yes,” we modify the original answers by flipping the last answer corresponding to Attempt to “No.” For example, if LLMs initially predict a post as {Yes, No, Yes} and the writer mentions moving on, we change the last “Yes” for Attempts to “No”. The refined answers then become {Yes, No, No}, and the post would be re-annotated as Ideation. Another approach we tried was to label posts as Indicator if the LLM initially predicted Attempt but the writer displayed positive signs within the post. However, this method did not produce favorable results on the validation dataset.
@@
We employ Qwen2-72B-Instruct  [^1] https://huggingface.co/Qwen/Qwen2-72B-Instruct to generate pseudo-label for the unlabeled posts, as it is one of the performant open-source models with an extended context window of up to $128K$ tokens [3]. We use greedy-decoding (i.e., choosing the highest probability output at each step of the sequence generation) and set the maximum length of the generated tokens to $1024$. We also experiment with other LLMs, such as LLaMA3 [4] and Mistral family models [37, 38], refer to Table VII and Section VI-A for further discussion.
@@
Combining models to generate training dataset. We use a combination of Qwen2-72B-Instruct, along with fine-tuned DepRoBERTa and Llama3-8B models on the $500$ labeled posts to generate pseudo-labels for unlabeled data, as depicted in Fig. 1(a), rightmost. Specifically, we only kept posts for which all three models predicted the same labels, and discarded the rest in case of disagreement. By doing so, we retain only high-confidence pseudo-labels, which allow us to reduce model-specific biases, resulting in a cleaner dataset for fine-tuning later on (refer to Table V in Section VI-A for more discussion). After labeling $1{,}500$ posts, we keep only $902$ posts with pseudo-labels. These posts, combined with the $500$ labeled dataset, form the training set of $1{,}402$ posts. Statistics for the four annotated suicide risk labels from the combined dataset are outlined in Table III.
@@ IV-C Ensemble Model for a robust and performant classifier
Our ensemble model consisting of five individual models: Qwen2-72B-Instruct via prompting, and four fine-tuned large language models (LLMs) - Llama3-8B (with two variants differing in hyperparameters, denoted as Llama3-8B 1 and Llama3-8B 2), Llama3.1-8B, and Gemma2-9B, as shown in Fig. 1(c). From $500$ labeled data, we assign the ensemble weight $\mathrm{w}_{i}$ to the models, with a weight of $2$ for Qwen2-72B-Instruct, and a weight of $1$ for the other four models. This means that the prediction of Qwen2-72B-Instruct counts twice, whereas all other models count once. The final prediction $\mathrm{p}$ is determined through a weighted majority voting over the five models as illustrated in Eq. 4. This ensemble approach combines the strengths of each individual model to create a more performant and robust classifier, as demonstrated and discussed in Section V and Section VI-B.
@@ VI Analysis and Discussion
VI-A Ablation Study: loss functions, model size for prompting, and noise reduction for pseudo-labels
@@
Figure 6: Correlation Between Model Predictions on 500 labeled data. The ensemble model demonstrates a stronger correlation with each model (last row and column), when compared to the pair of individual models.
@@
VI-B Model Evaluation and Performance
@@
VI-C Prediction Correlation and Token Distribution Analysis
Correlation Between Model Predictions. As shown in Fig. 4, the ensemble model demonstrates its robustness when evaluated on a new dataset. We compare the correlations between the predictions of the models, as illustrated in Fig. 6. Our observations indicate that the ensemble model exhibits a stronger correlation with each individual model compared to the correlations between pairs of individual models.
@@ VII Conclusion
However, our work has some limitations. The ensemble model, particularly Qwen2-72B-Instruct, requires a substantial amount of time for inference due to its large size, which poses challenges for real-time deployment. To mitigate this, smaller distillation models [44] could offer a more practical solution. Additionally, enhancing suicide detection using LLMs would benefit from better prompt engineering by domain experts. Breaking prompts into smaller, more specific questions can help direct LLMs to produce more relevant, accurate, and interpretable responses. Another promising approach is to collect more unlabeled data, leading to additional pseudo-labels for more effective fine-tuning of LLMs. Furthermore, incorporating visual data, such as images and videos from user posts, represents a promising direction for future research and may necessitate the use of multimodal LLMs.